OpenStudy (anonymous):

9^2n is less than equal to 1/6561

OpenStudy (anonymous):

n = -2 right ?

OpenStudy (anonymous):

yes, i finally got it!

OpenStudy (zzr0ck3r):

9^2n<= 1/6561 log_9(9^(2n)) = log_9(1/6561) 2n = log_9(1/6561) n = log_9(1/6561)/2 note: log_9() is log base 9